drawing pencils cost 8 cents each and colored pencils cost 11 cents each . Two dozen assorted pencils cost 2.16 dollars . how many colored pencils are There ?
Answer by tommyt3rd(5050)

1) x+y=24
2) ,08x+.11y=2.16
leads to
y=(2.16-0.08*24)/0.03=8
8-color pencils